OpenStudy (diamond123):

how do i reduce using the prime factorization method with fractions after multiplying ??????

OpenStudy (anonymous):

I answered this already.

OpenStudy (diamond123):

so i just cancel out the same numbers i see in both the denominator and numerator

OpenStudy (anonymous):

Yes, perhaps u can "see" them without actually factorizing but factorizing will always find anything common. If there is nothing in common, u cannot reduce any more.

OpenStudy (diamond123):

o oh okay so after i find the prime factorization of a fraction using multiplication i then look for the same numbers that are alike but what if the exponent is different

OpenStudy (anonymous):

Like 2^4 /2^3 ?

OpenStudy (diamond123):

yesssss

OpenStudy (anonymous):

2^4 = 2*2*2*2 2^3 = 2*2*2 Cancel 3 2's but there is a quicker way because they have already done the factorization for u. Just subtract exponents 4-3 = 1 -> 2^1 =2